OpenStudy (anonymous):

A regular hexagon has sides of 3 feet. What is the area of the hexagon?

OpenStudy (anonymous):

  the apothem= \[\frac{ 3\sqrt{3} }{ 2 }\] so, the area is: \[\frac{ 1 }{ 2 }(\frac{ 3\sqrt{3} }{ 2 })(18)\] \[A=\frac{ 27\sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

answer choices are 27 sqrt 3 , 18 sqrt 3, 9 sqrt 3, 13.5 sqrt 3

OpenStudy (anonymous):

\[27\div2=13.5\] so, the answer is \[13.5\sqrt{3}\] you just have to simplify it more.